# Self Iterative Label Refinement via Robust Unlabeled Learning

### v1

@@ 2. RQ2
    Can our method enhance performance on challenging tasks where even advanced LLMs (e.g., GPT-4o-based models) struggle?
@@ 3. RQ3
    Can our approach outperform existing LLM self-refinement strategies and recently introduced Reasoning models?
@@ 5.1 Experimental Setup
We use six binary classification datasets grouped into two categories based on their difficulty [^1] We evaluate difficulty based on a pilot experiment in which we trained the classification model in a standard supervised setting. Please refer to Table 1 for classification accuracy via supervised learning.. Table 1 reports the dataset statistics, and Table 2 provides examples for positive and negative cases (see Appendix A).
Easier Tasks (for RQ1): We evaluated our algorithm on three tasks: (i) Fake News (Ahmed et al., 2018), classifying news articles as fake or real, (ii) Saroco (Rogoz et al., 2021), Romanian satire detection dataset to assess effectiveness in a low-resource language, and (iii) Safety (Dai et al., 2024), dataset for SafeRLHF evaluating if the responses to questions is safe or dangerous, thereby assessing effectiveness for LLM’s post-training.
Harder Tasks (for RQ2 and RQ3): We evaluate our algorithm on three more challenging tasks: (i) the Corona Sentiment [^2] https://github.com/akshayjoshii/COVID19-Tweet-Sentiment-Analysis-and-EDA/tree/master, classifying social media post related to COVID-19 as positive or negative sentiment, (ii) the Green Patent [^3] https://huggingface.co/datasets/cwinkler/patents_green_plastics which involves identifying whether a patent abstract pertains to green plastics—requiring high expert knowledge, and (iii) Protein Structure (Blanchard et al., 2022), involving classification of proteins based on high or low COVID-19 binding affinity from molecular smiles strings.
For all experiments, we divided each dataset and randomly partitioned it into training, validation (for best epoch selection), and test splits in a 7:1:2 ratio.
@@
We train the classifier by appending an affine layer to the transformer’s final hidden state to yield a one-dimensional score. For fine-tuning efficiency, we employ QLora (Dettmers et al., 2023) with 4-bit quantization. At each iteration, we fine-tune the model using the pseudo-positive and pseudo-negative corpora with the AdamW optimizer (learning rate = $1.0\times 10^{-4}$, batch size = 16, and 3 epochs) and fix the robust UU learning hyperparameter $\lambda$ at -0.001 (see Equation (4)). At the end of each epoch, we compute the loss on a pseudo-labeled validation set and select the model with the lowest loss to re-label the entire dataset. Additional parameters are detailed in Table 3.
@@
Robust UU learning requires estimates of the positive proportions $\hat{\theta}_{p}$ and $\hat{\theta}_{n}$ for the pseudo-positive and pseudo-negative corpora, respectively. We compare two settings to obtain these parameters. In the first setting (Ours (Oracle)), we assume that the exact value of $\theta_{p}$ and $\theta_{n}$ are available [^4] The assumption that the precise class priors are known is not unrealistic given the effectiveness of existing class prior estimation methods (e.g., (Scott et al., 2013; Menon et al., 2015; Liu and Tao, 2016; Bekker and Davis, 2020)).. In the second setting (Ours (few-labeled)), only 50 examples from the training dataset are labeled, from which we compute $\hat{\theta}_{p}=p(y=1\mid\tilde{y}=1)$ and $\hat{\theta}_{n}=p(y=1\mid\tilde{y}=0)$ before proceeding with robust UU learning. Our experiments show that even a small amount of labeled data provides sufficiently accurate estimates.
@@
Figure 3: Classification accuracy over five iterations for three datasets. The solid lines represent the mean values, and the shaded areas show the mean $\pm$ standard deviation. Both variants of our approach, Ours (Oracle) and Ours (few-labeled) demonstrate steady improvements as the iteration increases. Notably, even in scenarios where the baselines fail to learn the classification task, our method continues to exhibit iterative performance gains. This robustness highlights the strength of our iterative refinement strategy, even under minimal supervision settings like 50 labeled examples.
@@ 5.2 Experiments on Easier NLP Tasks (RQ1)
We begin with three relatively simple tasks: Fake News, Saroco, and Safety. Given its relatively small parameter size and strong performance, we use Llama-3.2-1B-Instruct [^5] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct as the base model for the classifier.
@@
To generate the initial pseudo labels, we employ five open models, gemma-2-2b-it [^6] https://huggingface.co/google/gemma-2-2b-it, Llama-2-7b-chat-hf [^7] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Meta-Llama-3-8B [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B, Llama-3.2-1B-Instruct [^9] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct, and Llama-3.2-3B-Instruct [^10] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct to ensure reproducibility and to evaluate whether our self-refinement approach can operate robustly across a diverse set of models.
@@
We compare our approach against three baselines:
@@
    PN Learning treats the pseudo-labels as fully reliable, i.e., standard supervised training on potentially noisy labels.
@@
    PIE (Zhang et al., 2023) iterative label refinement method that temporarily accepts only high-confidence predictions as correct, training an ensemble-like model on these provisional labels and iteratively improving label accuracy.
@@
    CCP (Kutt et al., 2024) method leveraging an iterative contrastive learning pipeline that incrementally builds a set of highly reliable labels by learning robust class-specific features from a limited labeled dataset.
@@
Figure 3 shows the results for the easier tasks. Both Ours (Oracle) and Ours (few-labeled) steadily improve in accuracy, ultimately achieving high classification performance by the final iteration. Notably, despite using only 50 labeled examples, Ours (few-labeled) achieves performance comparable to Ours (Oracle), underscoring its potential even under minimal supervision.
For the Fake News task (using Meta-Llama-3-8B), PIE shows iterative performance gains, eventually matching the performance of Ours (Oracle) and outperforming standard PN Learning. This highlights the effectiveness of its ensemble strategy and confidence-based label refinement. However, for Saroco and Safety, where PN Learning struggles, PIE similarly performs poorly. This suggests that even with confidence-based filtering, limited initial classification accuracy hampers the effective elimination of noisy labels.
Similarly, while CCP demonstrates performance improvements with increasing iterations on the Fake News task, it completely fails on Saroco and Safety. CCP’s reliance on learning class-specific features from a small teacher dataset appears insufficient in scenarios where such features are inherently difficult to capture.
@@
5.3 Experiment on Harder Tasks (RQ2/RQ3)
We next evaluate our approach on three more challenging tasks: Corona Sentiment, Green Patent, and Protein Structure. To further explore the potential of our proposed method, we employed Llama-3.2-3B-Instruct for both Corona Sentiment and Green Patent and utilized bert-base-smiles [^11] https://huggingface.co/unikei/bert-base-smiles for Protein Structure. Initial pseudo labels are generated using the high-performance closed models GPT-4o-mini [^12] https://platform.openai.com/docs/models#gpt-4o-mini and GPT-4o [^13] https://platform.openai.com/docs/models#gpt-4o.
@@ 5.3 Experiment on Harder Tasks (RQ2/RQ3)
To compare the performance of self-refinement under a minimal human supervision setting, we adopt a self-refinement framework (Madaan et al., 2023; Kim et al., 2023; Chen et al., 2024b), where a response agent generates an initial answer and a feedback agent generates the feedback for this answer, thus iteratively refines answer (see Figures 7 and 8 for exact prompts). In this setup, we leverage high-performance closed models (GPT-4o-mini and GPT-4o) alongside the cost-effective reasoning model DeepSeek-R1 (DeepSeek-AI, 2025). For GPT-4o-mini and GPT-4o, initial annotations are generated consistently with our iterative robust UU learning (Ours) to ensure a fair comparison.
@@ Results:
In contrast, while the LLM-based self-refinement approach by GPT-4o-mini and GPT-4o shows a slight performance gain in the Corona Sentiment task, they suffer performance degradation in the Green Patent and Protein Structure tasks. Additionally, although DeepSeek-R1 starts with high annotation scores on all three datasets, it exhibits no performance gains in subsequent iterations. This suggests that even when employing a reasoning model, the benefits of self-refinement remain limited in scenarios where the LLM’s internal knowledge is insufficient. Therefore, relying solely on self-refinement in these cases may not lead to further performance gains.
Summary for RQ2 and RQ3: These findings confirm RQ2 by demonstrating our method enhances performance on challenging tasks through iterative refinement. They also confirm RQ3 by showing our approach outperforms existing LLM self-refinement strategies and reasoning models.
@@ 6 Conclusion
In this work, we introduced an iterative refinement pipeline that leverages robust UU learning to improve classification performance with minimal human supervision. Our approach mitigates noise in the initial annotations by decoupling pseudo-label refinement from the LLM’s internal judgments and by effectively harnessing the unlabeled corpus. Extensive experiments across diverse tasks—from fake news detection and low-resource language satire to challenging domains such as green patent identification and protein structure classification—demonstrate that our method consistently outperforms both the initial LLM annotations and existing self-refinement approaches.
Notably, our results show that even with as few as 50 labeled examples for fine-tuning, the few-labeled variant of our approach attains performance comparable to an oracle setting, highlighting its scalability and cost-effectiveness. This iterative framework not only enhances the quality of pseudo-labels over successive iterations but also proves robust in scenarios where conventional methods struggle, particularly in domains requiring specialized expertise or within AI for Science applications.
Future work may extend our framework to additional tasks such as LLM post-training and AI-driven scientific applications. Overall, our study demonstrates that iterative refinement combined with weakly supervised learning can significantly enhance LLM performance in low-resource, complex settings.
@@
Our approach has several limitations that warrant further exploration. Although our experiments confirm the method’s robustness under typical conditions, its performance might degrade when faced with extremely noisy pseudo-labels. For instance, the learning process becomes significantly more challenging when the positive prior for the pseudo-positive and pseudo-negative sets are nearly identical, resulting in an initial accuracy of around 0.5. In fact, as shown in Fig.3, in the Fake News task (using Gemma-2.2b), both our Oracle and few-labeled variants exhibit limited performance improvement, with an initial annotation accuracy of approximately 0.6. Conversely, Fig.1 shows that in the Protein Structure task (using GPT-4o), improvement is observed even when the initial annotation accuracy is around 0.55. Although these observations do not allow us to pinpoint a definitive threshold for ineffective initial annotations, they indicate that under conditions under extremely noisy annotation cases, the benefits of our iterative refinement framework will be limited.
@@
Acknowledgment
This research was supported by the JST Moonshot Research and Development Program JPMJMS2236-8.
Appendix A Dataset Details
A.1 Dataset Curation Details

### v2

@@ 2. RQ2
    Can our approach enhance performance on challenging tasks where even advanced LLMs (e.g., GPT-4o or DeepSeek-R1) struggle?
@@ 3. RQ3
    Can our method extend beyond classification tasks to generative tasks such as LLM alignment?
@@ 5.1 Experimental Setup
We use six binary classification datasets grouped into two categories based on their difficulty [^2] We evaluate difficulty based on a pilot experiment in which we trained the classification model in a standard supervised setting. Please refer to Table 2 for classification accuracy via supervised learning.. Table 2 reports the dataset statistics, and Table 3 provides examples for positive and negative cases (see Appendix B).
Easier Tasks (for RQ1): We evaluated our algorithm on three tasks: (i) Fake News [1], classifying news articles as fake or real, (ii) Saroco [44], Romanian satire detection dataset to assess effectiveness in a low-resource language, and (iii) Safety [12], dataset for SafeRLHF evaluating if the responses to questions is safe or dangerous, thereby assessing effectiveness for LLM’s post-training.
Harder Tasks (for RQ2): We evaluate our algorithm on three more challenging tasks: (i) the Corona Sentiment [^3] https://github.com/akshayjoshii/COVID19-Tweet-Sentiment-Analysis-and-EDA/tree/master, classifying social media post related to COVID-19 as positive or negative sentiment, (ii) the Green Patent [^4] https://huggingface.co/datasets/cwinkler/patents_green_plastics which involves identifying whether a patent abstract pertains to green plastics, requiring high expert knowledge, and (iii) Protein Structure [7], involving classification of proteins based on high or low COVID-19 binding affinity from molecular smiles strings.
For all experiments, we randomly partitioned each dataset into training, validation (for best epoch selection), and test splits in a 7:1:2 ratio.
@@
We train the classifier by appending an affine layer to the transformer’s final hidden state to yield a one-dimensional score. For fine-tuning efficiency, we employ QLora [14] with 4-bit quantization. At each iteration, we fine-tune the model using the pseudo-positive and pseudo-negative corpora with the AdamW optimizer (learning rate = $1.0\times 10^{-4}$, batch size = 16, and 3 epochs) and fix the robust UU learning hyperparameter $\lambda$ at -0.001. At the end of each epoch, we compute the loss on a pseudo-labeled validation set and select the model with the lowest loss to re-label the entire dataset. Additional parameters are detailed in Table 4.
@@
Robust UU learning requires estimates of the positive proportions $\hat{\theta}_{p}$ and $\hat{\theta}_{n}$ for the pseudo-positive and pseudo-negative corpora, respectively. We compare two settings to obtain these parameters. In the first setting (Ours (Oracle)), we use the exact values of $\theta_{p}$ and $\theta_{n}$, which represents a theoretical upper bound on performance [^5] The assumption that the precise class priors are known is not unrealistic given the effectiveness of existing class prior estimation methods (e.g., [5, 33, 38, 47]).. In the second setting (Ours (few-labeled)), we assume a practical scenario with only a small subset of labeled data (50 or 100 examples). From these examples, we estimate the class priors by computing $\hat{\theta}_{p}=p(y=1\mid\tilde{y}=1)$ and $\hat{\theta}_{n}=p(y=1\mid\tilde{y}=0)$ before applying robust UU learning. Our experiments show that even a small amount of labeled data provides sufficiently accurate estimates.
@@
Figure 2: Classification accuracy over five iterations for three datasets. The solid lines represent the mean values, and the shaded areas show the mean $\pm$ standard deviation. Both variants of our approach, Ours (Oracle) and Ours (few-labeled), demonstrate steady improvements as the iteration increases. Notably, even in scenarios where the baselines fail to learn the classification task, our method continues to exhibit iterative performance gains. This robustness highlights the strength of our iterative refinement strategy, even under minimal supervision settings like 50 labeled examples. Detailed numerical results are provided in Appendix D.
@@ 5.2 Experiments on Easier NLP Tasks (RQ1)
We start with three relatively simple tasks: Fake News, Saroco, and Safety. For the base model of the classifier, we use Llama-3.2-1B-Instruct [^6] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct, a compact language model that offers strong performance with only 1 billion parameters.
@@
To generate the initial pseudo-labels, we employ five open models of varying sizes and model families: gemma-2-2b-it (2B) [^7] https://huggingface.co/google/gemma-2-2b-it, Llama-2-7b-chat-hf (7B) [^8] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Meta-Llama-3-8B (8B) [^9] https://huggingface.co/meta-llama/Meta-Llama-3-8B, Llama-3.2-1B-Instruct (1B) [^10] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct, and Llama-3.2-3B-Instruct (3B) [^11] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct. This diverse selection helps ensure reproducibility and allows us to evaluate whether our self-refinement approach operates robustly across different model families and scales.
@@
Table 1: Overview of experimental baselines and their categories.
Method Category
PN Vanilla Supervised
PIE [66] Weakly Supervised
CCP [26] Semi-Supervised
$U$ Ablation (No Robust)
Ours (Oracle) Ablation (Ceiling)
Ours (50-labeled) Proposed
Ours (100-labeled) Proposed
We compare our approach against three baseline methods that cover different learning paradigms:
@@
    PN: Standard supervised training on pseudo-labels, treating them as fully reliable.
@@
    PIE [66]: Iterative method that accepts high-confidence predictions as correct labels, progressively training ensemble models on these provisional labels to iteratively improve classification accuracy.
@@
    CCP [26]: Semi-supervised method using iterative contrastive learning to construct reliable pseudo-labels by learning robust class-specific feature representations from a limited set of labeled data.
We also conducted an ablation study comparing standard UU learning (2) against our robust UU learning with various prior estimation methods. Table 1 summarizes all methods.
@@
Figure 2 shows the results for the easier tasks. All three variants, Ours (Oracle), Ours (50-labeled), and Ours (100-labeled), show steady accuracy gains and reach high performance by the final iteration. Notably, the two limited-label settings quickly converge to the Oracle upper bound, yielding comparable final accuracies despite relying on only 50 or 100 labels. This outcome aligns with the robustness of UU learning to class-prior estimation errors [35], enabling strong performance under severe label scarcity and underscoring the method’s practicality in minimal-supervision scenarios.
For the Fake News task (using Meta-Llama-3-8B), PIE shows iterative performance gains, eventually matching the performance of Ours (Oracle) and outperforming PN Learning. This highlights the effectiveness of its ensemble strategy and confidence-based label refinement. However, for Saroco and Safety, where PN Learning struggles, PIE similarly performs poorly, suggesting that even with confidence-based filtering, limited initial classification accuracy hampers noisy labels elimination.
Similarly, while CCP shows performance gains with increasing iterations on the Fake News task, it completely fails on Saroco and Safety. CCP’s reliance on learning class-specific features from a small teacher dataset appears insufficient where such features are inherently difficult to capture.
The UU baseline also improves with each iteration. However, its peak accuracy consistently falls short of oracle and few-labeled variants of our approach, confirming the advantage of our robust correction for noisy pseudo-labels.
While our method demonstrates strong performance across diverse tasks, its gains can be limited when the initial pseudo-labels are highly noisy, causing the class priors of the pseudo-positive and pseudo-negative sets to be nearly identical. As shown in Figure 2, the Fake News task (using Gemma-2.2b) exhibits restricted improvements, starting with an initial annotation accuracy of 0.591. Conversely, the Saroco task (Llama-3.2-1B) achieves clear iterative gains despite a similarly low initial accuracy (0.576). This suggests that while the initial noise level influences performance, it does not solely determine the success of the refinement process.
@@
5.3 Experiments on Harder Tasks (RQ2)
We next evaluate our approach on three more challenging tasks: Corona Sentiment, Green Patent, and Protein Structure. To further explore the potential of our method, we employed Llama-3.2-3B-Instruct for both Corona Sentiment and Green Patent and utilized bert-base-smiles [^12] https://huggingface.co/unikei/bert-base-smiles for Protein Structure. Initial pseudo labels are generated using the high-performance closed models GPT-4o-mini [^13] https://platform.openai.com/docs/models#gpt-4o-mini and GPT-4o [^14] https://platform.openai.com/docs/models#gpt-4o.
@@ 5.3 Experiments on Harder Tasks (RQ2)
To compare the performance of self-refinement under a minimal human supervision setting, we adopt a self-refinement framework [11, 23, 37], where a response agent generates an initial answer and a feedback agent generates the feedback for this answer, thus iteratively refines answer (see Figures 6 and 7 for exact prompts). In this setup, we leverage high-performance closed models (GPT-4o-mini and GPT-4o) alongside the cost-effective reasoning model DeepSeek-R1 [13]. For GPT-4o-mini and GPT-4o, initial annotations are generated consistently with our iterative robust UU learning (Ours) to ensure a fair comparison.
Figure 3: Classification accuracy curves over five iterations on three challenging datasets. Ours (Oracle) uses the exact class prior for UU learning, while Ours (few-labeled) estimates these priors from only 50 labeled examples. Our method consistently improves accuracy and outperforms both LLM self-refinement by GPT-4o series and advanced reasoning model DeepSeek-R1. Detailed numerical results are provided in Appendix D.
@@ Results:
In contrast, the LLM-based self-refinement approach by GPT-4o-mini and GPT-4o shows only a slight performance gain in the Corona Sentiment task and suffers from performance degradation in the Green Patent and Protein Structure tasks, where accuracy actually worsens over iterations. Similarly, although DeepSeek-R1 starts with high annotation scores on all three datasets, its performance plateaus, exhibiting no gains in subsequent iterations. These results suggest that even when employing a sophisticated reasoning model, the benefits of self-refinement are limited when the LLM’s internal knowledge is insufficient to correctly evaluate and revise its own outputs. Therefore, relying solely on self-refinement in these challenging domains may not lead to further performance gains and can even be counterproductive.
Summary for RQ2: These experimental findings answer RQ2: While self-refinement with advanced LLMs like GPT-4o and DeepSeek-R1 fails to improve or even degrades performance on challenging tasks, our method consistently enhances accuracy through iterative refinement, ultimately outperforming these strong baselines.
5.4 LLM Alignment on Safety (RQ3)
Setup:
We reuse the Safety dataset as a generative benchmark for safety alignment, employing a RLHF approach. The reward model (RM) was a classifier trained by our robust UU pipeline: we use llama-3.2-1B-Instruct, which provides initial pseudo-labels and as the base model for the classifier. This refined classifier then functioned as the RM.
As the base policy for RLHF, we applied supervised fine-tuning (SFT) to llama-3.2-1B on the Alpaca open-source dataset [54]. By intentionally using models from the same llama-3.2-1B series for the base model, pseudo-label generation, and the reward model, we aimed to evaluate the potential of self-refinement in generative tasks. See the Appendix C.2 for experimental details and our training and evaluation procedure following [12].
Baselines:
We compare four systems: (i) SFT, (ii) Vanilla RLAIF that employs an RM trained only with PN learning, (iii) Ours (Oracle) that uses the true priors, and (iv) Ours (50-labeled) that estimates class priors from 50 labeled examples. All experiments are repeated with three random seeds; we aggregate reward distributions over the test prompts.
Figure 4: Reward distribution of generated answers on the Safety dataset after alignment. Both variants of our method, Ours (Oracle) and Ours (50-labeled), shift the distribution toward higher (safer) rewards compared with the SFT baseline and the Vanilla RLAIF. Legend values denote the mean reward across three random seeds.
Results:
Figure 4 shows that while the SFT policy still produces a noticeable fraction of negative-reward (unsafe) outputs, both our oracle and few-labeled variants dramatically reduce this undesirable left tail, concentrating the probability distribution within the safe region. In contrast, Vanilla RLAIF fails to achieve meaningful improvement, performing worse than the baseline SFT, which emphasizes that noisy reward signals severely hamper RL optimization. The robust reward modeling provided by our UU-refinement strategy delivers a stable, noise-resistant learning signal, effectively guiding the generative model toward safer and more desirable behavior.
Summary for RQ3:
These findings answer RQ3: the improved classification performance achieved through our iterative robust UU learning strategy successfully translates to the more challenging generative alignment task. This result confirms that our approach provides a practical and highly effective self-refinement mechanism, beneficial not only in classification but also in complex generative settings.
@@ 6 Conclusion
We proposed an iterative refinement pipeline leveraging robust UU learning with minimal supervision, which consistently outperforms raw LLM annotations and existing self-refinement approaches across six classification benchmarks. Notably, our method achieves performance comparable to oracle-prior settings using only 50 labeled examples. It also demonstrates significant effectiveness in generative tasks, successfully enabling RLAIF on the Safety dataset, where naive approaches previously failed. Our approach offers a simple yet powerful means to enhance LLM performance, particularly in low-resource and complex domains, with substantial potential for annotation-intensive applications such as advanced LLM self-refinement and AI for Science.
Limitations and Future Work.
Although our method is robust, its effectiveness can be limited by extreme initial pseudo-label noise. Our findings indicate that performance depends not only on the quality of initial labels but also significantly on other contextual factors. A promising direction for future work involves explicitly modeling instance-level classifiability, acknowledging the variable difficulty between task formulations, such as chat-style judgments by language models versus predictions from trained classifiers, an example easy for one may prove challenging for the other. Utilizing such instance-specific difficulty metrics to inform weighting, sampling, and curriculum strategies could prioritize reliable examples while incorporating challenging yet informative cases. Additionally, augmenting classifiers with auxiliary information, such as rationales, retrieved contexts, or lightweight metadata, could further improve classification accuracy for ambiguous texts.
Acknowledgments
This work has been supported by the JST Moonshot Research and Development Program JPMJMS2236-8.
Appendix A Limitations and Risks
@@
Our approach has several limitations that warrant further exploration. As we discussed in Section 5.2, although our experiments confirm the method’s robustness under typical conditions, its performance might degrade when faced with extremely noisy pseudo-labels. For instance, the learning process becomes significantly more challenging when the positive prior for the pseudo-positive and pseudo-negative sets are nearly identical, resulting in an initial accuracy of around 0.5. In fact, as shown in Figure 2, in the Fake News task (using Gemma-2.2b), both our Oracle and few-labeled variants exhibit limited performance improvement, with an initial annotation accuracy of approximately 0.6. Conversely, Figure 3 shows that in the Protein Structure task (using GPT-4o), improvement is observed even when the initial annotation accuracy is around 0.55. Although these observations do not allow us to pinpoint a definitive threshold for ineffective initial annotations, they indicate that under conditions under extremely noisy annotation cases, the benefits of our iterative refinement framework will be limited.
@@
Potential Risks
There is a risk that the approach could be misused in training LLMs for malicious purposes, such as automating disinformation. Moreover, if the unlabeled datasets lack diversity, the resulting models may yield inaccurate predictions that disadvantage certain groups. Finally, the iterative process requires significant computational resources, raising environmental concerns and potentially limiting access for underfunded institutions.
Appendix B Dataset Details
B.1 Dataset Curation Details